OpenStudy (anonymous):

Solve using substitution: 9y - x - 8 = 0 -3x + 7y = -16

OpenStudy (anonymous):

(-8, 0) (1, 1) (0, 8/9) (10, 2)

OpenStudy (triciaal):

from equation 1 write the expression for the variable x then substitute this in the 2nd equation

OpenStudy (triciaal):

this gives one equation with one variable; solve for y

OpenStudy (triciaal):

now put this value in the expression to find x

OpenStudy (triciaal):

do you need more help or can you follow these directions to get D?

OpenStudy (triciaal):

It is also a good practice to check your answer in the original

OpenStudy (hockeychick23):

(10,2) Heres your check by plugging it into the equation: 9y-x-8= 0 9(2) - 10 - 8 = 0 18-10-8=0 -3x + 7y = -16 -3(10)+7(2)=-16 -30+14=-16